# Revisiting Differentially Private Hyper-parameter Tuning

### v2

@@ Abstract
We study the application of differential privacy in hyper-parameter tuning, a crucial process in machine learning involving selecting the best hyper-parameter from several candidates. Unlike many private learning algorithms, including the prevalent DP-SGD, the privacy implications of tuning remain insufficiently understood or often totally ignored. Recent works propose a generic private selection solution for the tuning process, yet a fundamental question persists: is this privacy bound tight?
This paper provides an in-depth examination of this question. Initially, we provide studies affirming the current privacy analysis for private selection is indeed tight in general. However, when we specifically study the hyper-parameter tuning problem in a white-box setting, such tightness no longer holds. This is first demonstrated by applying privacy audit on the tuning process. Our findings underscore a substantial gap between current theoretical privacy bound and the empirical bound derived even under strong audit setups.
This gap motivates our subsequent investigations. Our further study provides improved privacy results for private hyper-parameter tuning due to its distinct properties. Our results demonstrate broader applicability compared to prior analyses, which are limited to specific parameter configurations.
Index Terms:
Hyper-parameter Tuning; Privacy Audit; Private Selection; Differential Privacy;
1 Introduction
Differential Privacy (DP) [18] stands as the prevailing standard for ensuring privacy in contemporary machine learning. A ubiquitous technique employed to ensure DP across a diverse array of machine learning tasks is differentially private stochastic gradient descent (DP-SGD, a.k.a., noisy-SGD) [6, 45, 2]. The ongoing refinement of DP-SGD’s privacy analysis [2, 4, 33, 35], referred to as “privacy accounting” or “privacy budgeting”, has contributed to a profound comprehension of the mechanism.
@@ 1 Introduction
Naively, one can bound the privacy loss for the tuning operation by the composition theorem. If we run the private base algorithm $k$ times with different hyper-parameters, the total privacy cost deteriorates at most linearly with $k$ (or $\mathcal{O}(\sqrt{k})$ if the base algorithm satisfies approximate DP and we use advanced composition theorem [20], which is nearly optimal [24]). However, these bounds are still far from satisfactory as $k$ is usually large in practice. Perhaps due to this limitation, it still remains common to exhaustively tune a private algorithm to achieve strong performance but only consider the privacy cost for a single run [15, 56, 54, 43].
@@
Thanks to the contributions of Liu and Talwar [30], hyper-parameter tuning now enjoys significantly better privacy bound than naively applying composition theorem. To briefly describe their findings, if we run a private base algorithm a random number of times (possibly with different hyper-parameters) and only output the best single run, the privacy cost only deteriorates by a constant multiplicative factor [30]. For example, if the base algorithm is $(\varepsilon,0)$-DP, then the whole tuning process is $(3\varepsilon,0)$-DP if the running number follows a geometric distribution [30]. This is much better than the $(k\varepsilon,0)$-DP bound under fixed $k$ times of running.
@@
Motivations. Although hyper-parameters can be tuned with rigorous privacy, whether the enhanced privacy analysis [30, 42] is tight or not is still an open problem. We note that results provided by [30] and [42] for certain setups show that the privacy cost still increases substantially (e.g., tripling compared to the base algorithm’s privacy cost as shown above); however, it also seems plausible that only revealing the best single run should not consume that much privacy budget. This prompts a natural question: Does hyper-parameter tuning truly consume a noticeably larger privacy budget than the base algorithm, as indicated, for instance, by a factor of roughly three [30]?
A negative answer would meaningfully imply that a more nuanced privacy-utility trade-off can be achieved through analysis alone. Conversely, a positive answer would indicate significant improvement in analysis is impossible. Overall, in contrast to the well-established understanding of privacy deterioration due to composition, investigating such a problem contributes to a deeper understanding of how privacy degrades due to an alternative factor: selection.
@@
This work. We answer the posed question with both positive and negative answers. In the affirmative, our constructed example demonstrates that the current generic privacy bound provided in [42] for private selection is indeed tight. Still, the result only holds in the worst case. Conversely, in the negative, we uncover a more favorable privacy bound given the base algorithm is DP-SGD specifically. Our contributions are as follows.
@@
1) Validating tightness of generic privacy bound for private selection (Section 3.2). We first provide a private selection instance where we observe only a negligible gap between the true privacy cost (with our detailed analysis) and the cost predicted by the current privacy bound [42]. However, when we study the private hyper-parameter tuning problem, where the base algorithm is DP-SGD, such tightness no longer holds. This finding is related to our other two contributions.
@@
2) Empirical investigation on the privacy leakage of hyper-parameter tuning (Section 4). We first take empirical approaches to investigate how much privacy is leaked when performing hyper-parameter tuning. This is done via the privacy audit technique [39, 24, 23], an interactive protocol used to empirically measure the privacy of some mechanisms. In contrast, unlike all previous privacy auditing work, which focuses on the privacy of the base algorithm (e.g., DP-SGD), auditing the tuning procedure is a fresh problem that requires new formulation and insight. Specifically, the score function, used to select the “best”, is the new factor that needs to be settled.
@@
We formulate various privacy threat models tailored for hyper-parameter tuning, where the weakest one corresponds to the most practical scenario and the strongest one corresponds to the worst case. Results under the weakest provide evidence that the tuning process hardly incurs additional privacy costs beyond the base algorithm. Notably, even the empirical privacy bound derived from the strongest adversary still exhibits a substantial gap from the generic privacy bound proposed by [42]. This gap motivates us to derive better theoretical results in the remaining sections.
@@
3) Improved privacy results (Sections 5.2, 5.3 and 6). We find that tuning a DP-SGD protocol does enjoy a better privacy result, as demonstrated by our subsequent study on deriving an improved privacy result. The pivotal aspect driving this improvement lies in representing the privacy of the base algorithm with finer resolution, and DP-SGD does have a distinctive characterization. This is done within the $f$-DP framework [17], deviating from the well-known $(\varepsilon,\delta)$-DP [18] or RDP [33]. We show that our results are tight in a general sense. Our results are also more generalizable, contrasting to previous work [30, 42], where they only easily apply to a limited range of parameter setups.
Subsequent to our improved results is a further experimental evaluation. We aim to compare our improved privacy result with the empirical privacy lower bound derived under an idealized audit setup. Notably, there is still a gap in between. This finding is examined in detail, revealing that the score function, a new factor in auditing hyper-parameter tuning, is a key determinant influencing audit performance. To give a concise summary of our main contributions:
  • We confirm the general tightness of the current generic privacy bound for private selection.
  • We empirically evaluate the privacy leakage due to hyper-parameter tuning.
@@
  • We provide improved privacy results for private hyper-parameter tuning.
2 Background
2.1 Differential Privacy (DP)
@@
Differentially private algorithms are resilient to post-processing, and the execution of multiple DP algorithms sequentially, known as composition, also maintains DP. Rényi DP (RDP), a DP relaxation shown in the following, often serves as a tight analytical tool to assess the privacy cost under composition.
@@
Differentially private stochastic gradient descent (DP-SGD) [6, 45, 2]. A machine learning model denoted as $f_{w}$ often represents a neural network with trainable parameters $w$. Specifically tailored for classification tasks in this study, $f_{w}$ may take image data as input and output corresponding labels. Stochastic Gradient Descent (SGD) [28] is the method for updating $w$ iteratively. Hyper-parameters, such as the updating step size (a.k.a. learning rate), often need to be tuned to gain optimized performance.
As the private version of SGD, DP-SGD ❶ computes the per-sample gradient for each sub-sample data, ❷ clips each gradient to have bounded $l_{2}$ norm, and ❸ adds Gaussian noise (Gaussian mechanism). The resultant private gradient $p_{i}$ is used to update parameter $w$, and $i$-th (up to $N$-th) update can be expressed as:
@@
where $\mathbf{lr}$ is the learning rate and data batch $\mathbf{B}$ contains the sub-sampled data (e.g., $x$ may represent an image, and $y$ is its label) where the sample ratio is $\tau$. The function $\operatorname{CLP}_{C}(u)=u\cdot\min(1,\frac{C}{\|u\|_{2}})$ where the clipping threshold $C$ is a hyper-parameter. $R_{i}$ is calibrated isotropic Gaussian noise sampled from $\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ where $d$ is the dimension or number of trainable parameters and $\sigma$ is the noise multiplier. $\ell(;)$ is the loss function that represents the neural network $f_{w}$ and some loss metric (e.g. cross-entropy loss). With $R_{i}$ and $\operatorname{CLP}_{C}$ operation eliminated, we recover SGD. There are notable variants of DP-SGD, such as DP-Adam [48]; they share the same privacy analysis as DP-SGD because of the post-processing property of DP.
2.2 Privacy Auditing
@@
$\mathrm{FP}+\mathrm{e}^{\varepsilon}\mathrm{FN}\geq 1-\delta,\text{\quad}\mathrm{FN}+\mathrm{e}^{\varepsilon}\mathrm{FP}\geq 1-\delta$ (3)
@@
2:$b_{\text{truth}}\leftarrow\{0,1\}$triangleright Trainer flips a fair coin
3:$\hat{X}\leftarrow X$ if $b_{\text{truth}}=0$, $\hat{X}\leftarrow X^{\prime}$ otherwise
4:Run $\mathcal{P}(\hat{X})$triangleright Trainer runs the private protocol
5:$b_{\text{guess}}\leftarrow\{0,1\}$triangleright Adversary makes a guess based on $\mathcal{P}(\hat{X})$
6:$(b_{\text{truth}},b_{\text{guess}})$
@@
Privacy audit. Privacy audit aims to find a lower bound of the privacy cost for a private protocol $\mathcal{P}$ based on the hypothesis testing interpretation of DP as shown above. This is usually done via simulating the interactive game-based protocol described in Algorithm 1. Such a simulation is typically repeated many times, resulting in many pairs of $(b_{\text{truth}},b_{\text{guess}})$. Then, the $\mathrm{FP}$ and $\mathrm{FN}$ for adversary’s guessing are computed by Clopper-Pearson method [11], which means that lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ computed by Equation (4) is with a confidence specification. If the adversary can make very accurate guesses and derive a lower bound higher than some claimed privacy parameter, it suggests $\mathcal{P}$ is not private as claimed.
@@
Related work on privacy audit. In privacy-preserving machine learning, privacy audit mainly serves a different goal from that of certain earlier studies [52, 16, 8, 7] on detecting privacy violation in general query-answering applications. Previous work on privacy audit in machine learning mainly targets auditing the DP-SGD protocol to assess its theoretical versus practical privacy [39, 24, 23]. Additional studies [47, 37, 31, 57] concentrate on enhancing the strength of audits on DP-SGD (yielding stronger/larger-value lower bound) or improving the efficiency (incurring fewer simulation overheads). Drawing a parallel to the action of guessing whether a data point was included or not, privacy audit may also be linked to membership inference attack (MIA) [38, 44], but diverges from MIA in terms of both their goals and focus. There are also recent works on auditing prediction [9] and synthetic data generation [3], which differ from our auditing experiments.
2.3 Private Hyper-parameter Tuning
@@
Problem Formulation. We formulate the private hyperparameter tuning problem aligning with [30, 42]. Let $\Omega=\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{m}\}$ be a collection of DP-SGD algorithms. These correspond to $m$ possible hyper-parameter configurations. We have $\mathcal{M}_{i}:\mathcal{X}\rightarrow\mathcal{Y}$ for $i\in[m]$, and all of these algorithms satisfy the same privacy parameter, i.e., they are all $(\varepsilon,\delta)$-DP for the same $\varepsilon,\delta$. The practitioner can freely determine the size of $\Omega$.
@@
The goal is to return an algorithm element (including its execution) of $\Omega$ such that the output of such algorithm has (approximately) the maximum score as specified by some score function $g:\mathcal{Y}\rightarrow\mathbb{R}$. The score function $g$ usually serves a utility purpose (e.g., $g$ could evaluate the validation loss on a held-out dataset). The selection must be performed in a differentially private manner. The general private selection problem corresponds to the cases where $\Omega$ contains some arbitrary differentially private algorithms.
@@
Related work on private hyper-parameter tuning. Well-established algorithms like the sparse vector technique [19] and exponential mechanism [32] may potentially be leveraged to the tuning problem; however, they assume a low sensitivity in the metric defining the “best”, a condition not always applicable. Some earlier work [10] also suffers from the same issue. Papernot et al. [42] and Liu [30] have provided generic private selection approaches circumventing such challenges. Mohapatra et al. [36] study privacy issues in adaptive hyper-parameter tuning under DP, which is different from the non-adaptive tuning problem considered in this work. There is related work [25] that builds upon [42]; therefore, what we understand about the generic approach in this study also applies to [25].
@@
Focus of this paper. Our first focus is to leverage and formulate specific privacy audit instantiation to understand how privacy deteriorates due to selection, diverging from all previous privacy audit work on privacy deteriorating due to composition. Furthermore, we also study improving privacy results specifically for a white-box application: the hyper-parameter tuning problem, pre-conditioned on the base algorithm is DP-SGD.
Algorithm 2 Private Selection Protocol $\mathcal{H}$

### v3

@@ Abstract
We investigate the application of differential privacy in hyper-parameter tuning, a process involving selecting the best run from several candidates. Unlike many private learning algorithms, including the prevalent DP-SGD, the privacy implications of selecting the best are often overlooked. While recent works propose a generic private selection solution for the tuning process, an open question persists: is such privacy upper bound tight?
This paper provides both empirical and theoretical examinations of this question. Initially, we provide studies affirming the current privacy analysis for private selection is indeed tight in general. However, when we specifically study the hyper-parameter tuning problem in a white-box setting, such tightness no longer holds. This is first demonstrated by applying privacy audit on the tuning process. Our findings underscore a substantial gap between the current theoretical privacy bound and the empirical privacy leakage derived even under strong audit setups.
This gap motivates our subsequent theoretical investigations, which provide improved privacy upper bound for private hyper-parameter tuning due to its distinct properties. Our improved bound leads to better utility. Our analysis also demonstrates broader applicability compared to prior analyses, which are limited to specific parameter configurations. Overall, we contribute to a better understanding of how privacy degrades due to selection  [^1] To appear in Network and Distributed System Security (NDSS) Symposium 2026.
I Introduction
Differential Privacy (DP) [18] stands as the prevailing standard for ensuring privacy in contemporary machine learning. A ubiquitous technique employed to ensure DP across a diverse array of machine learning tasks is differentially private stochastic gradient descent (DP-SGD, a.k.a., noisy-SGD) [5, 45, 2].
@@ I Introduction
Naively, one can bound the privacy loss for the tuning operation by the composition theorem. If we run the private base algorithm $k$ times with different hyper-parameters, the total privacy cost deteriorates at most linearly with $k$ (or $\mathcal{O}(\sqrt{k})$ if the base algorithm is approximate DP and we use advanced composition theorem [20], which is nearly optimal [24]). However, these bounds are still far from satisfactory as $k$ is usually large in practice. Perhaps due to this limitation, it still remains common to exhaustively tune a private algorithm to achieve strong performance but only consider the privacy cost for a single run [14, 56, 54, 43].
@@
Thanks to Liu and Talwar [30], hyper-parameter tuning now enjoys significantly better privacy bound than naively applying composition theorem. To briefly describe their findings, if we run a private base algorithm a random number of times (possibly with different hyper-parameters) and only output the best single run, the privacy cost only deteriorates by a constant multiplicative factor [30]. For example, if the base algorithm is $(\varepsilon,0)$-DP, then the whole tuning process is $(3\varepsilon,0)$-DP if the running number follows a geometric distribution [30]. This is much better than the $(k\varepsilon,0)$-DP bound under fixed $k$ times of running.
@@
Motivations. Hyper-parameters can be tuned with formal privacy guarantees. But it remains unclear whether the existing privacy analyses [30, 42] are tight.
Some results suggest the cost of tuning is high. For example, [30] shows the privacy budget can increase by a factor of three in certain setups. Still, this seems counterintuitive. It seems plausible that only revealing the best single run should not consume that much privacy budget. This leads to our core question: Does hyperparameter tuning actually consume significantly more privacy than the base algorithm?
If the answer is positive, then further significant improvements in the analysis is impossible. If negative, it is still valuable to pursue tighter analysis, as tighter bounds would allow more hyperparameter trials under the same budget, directly improving utility in private model selection.
@@
This work. We answer the posed question with both positive and negative answers. In the affirmative, our constructed example demonstrates that the current generic privacy bound provided in [42] for private selection is indeed tight. Still, the result only holds in the worst case. Conversely, in the negative, we uncover a more favorable privacy bound given the base algorithm is specific DP-SGD. We aim to understand how selection leaks privacy, in contrast to the well-established understanding of privacy deterioration due to composition. Our contributions are as follows.
@@
1) Validating tightness of generic privacy bound for private selection (Section III-B). We first provide a private selection instance where we observe only a negligible gap between the true privacy cost and the cost predicted by the current privacy bound [42]. Such results meaningfully show that significant improvement in privacy upper bound is impossible in general. However, when we study the private hyper-parameter tuning problem, where the base algorithm is DP-SGD, we enjoy better upper bounds. This finding is related to our other two contributions.
@@
2) Empirical investigation on how hyper-parameter tuning (selection) leaks privacy (Section IV). We first take empirical approaches to investigate how much privacy is leaked when performing hyper-parameter tuning. This is done via the privacy audit technique [39, 24, 23], an interactive protocol used to empirically measure the privacy of some mechanisms.
In contrast, unlike all previous privacy auditing work, which focuses on the privacy of the base algorithm (e.g., DP-SGD), auditing the tuning procedure is a fresh problem that requires new formulation and insight. Specifically, the score function used to select the “best” is the new factor that must be settled.
@@
We formulate various privacy threat models tailored for hyper-parameter tuning, where the weakest one corresponds to the most practical scenario and the strongest one corresponds to the worst case. Results under the weakest provide evidence that the tuning process hardly incurs additional privacy costs beyond the base algorithm. Moreover, even the empirical privacy bound (lower bound) derived from the strongest adversary still exhibits a substantial gap from the generic privacy bound (upper bound) proposed by [42]. In contrast, previously, the gap (between privacy lower bound and upper bound) is essentially closed in the audit on DP-SGD’s privacy [39]. Why are different audit results seen in auditing hyper-parameter tuning? This motivates our subsequent theoretical investigations.
@@
3) Improved theoretical privacy results (Sections V-B, V-C and VI). Our subsequent study shows that tuning DP-SGD does enjoy a better privacy result. The pivotal aspect driving this improvement lies in representing the privacy of the base algorithm with finer resolution, and DP-SGD does have a distinctive characterization. This is done within the $f$-DP framework [17], deviating from the well-known $(\varepsilon,\delta)$-DP [18] or RDP [33].
@@
Our improved result directly benefits differentially private hyper-parameter tuning:
it allows us to test substantially more (in expectation) hyper-parameters without increasing privacy budget, which translates to improved utility.
Our results are generalizable, contrasting to previous work [30, 42], which remains unknown how to adapt to general parameter setups.
Subsequent to our improved results is a further empirical evaluation: comparing our improved theoretical privacy result with the empirical privacy lower bound derived under an idealized audit setup. Interestingly, there is still a gap in between. This finding is examined in detail, revealing that the score function, a new factor in auditing hyper-parameter tuning, is a key determinant influencing audit performance. Consequently, this also prompts an exciting and essential open problem in the future: how to close such a gap.
II Background
II-A Differential Privacy (DP)
@@
We introduce Rényi DP (RDP), a DP relaxation shown in the following, often serves as a tight analytical tool to assess the privacy cost under composition.
@@
Differentially private stochastic gradient descent (DP-SGD) [5, 45, 2]. We use a machine learning model $f_{w}$, typically a neural network with trainable parameters $w$. In our classification setting, $f_{w}$ maps inputs (e.g., images) to labels. Parameters are updated using Stochastic Gradient Descent (SGD) [28], where hyper-parameters like learning rate must be tuned for good performance.
DP-SGD is the private version of SGD. It follows three steps: 1) compute per-sample gradients; 2) clip each to have bounded $\ell_{2}$ norm; 3) add Gaussian noise.
The private gradient $p_{i}$ is then used to update $w$ as:
@@
Here, $\mathbf{B}$ is the sampled batch (with ratio $\tau$), $\mathbf{lr}$ is the learning rate, and $\ell$ is the loss (e.g., cross-entropy). Clipping is defined as $\operatorname{CLP}_{C}(u)=u\cdot\min(1,\frac{C}{|u|_{2}})$, where $C$ is a clipping threshold. Noise $R_{i}$ is sampled from $\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$, where $\sigma$ is the noise multiplier and $d$ is the number of parameters. Removing clipping and noise recovers (mini-batch) SGD. Variants like DP-Adam [48] follow the same privacy analysis by the post-processing property of DP.
II-B Privacy Audit
@@
[…]
@@
[…]
@@
Privacy audit. Privacy audit aims to find a lower bound of the privacy cost for a private protocol $\mathcal{P}$ based on the hypothesis testing interpretation of DP as shown above. This is usually done via simulating the interactive game-based protocol described in Algorithm 1. Such a simulation is typically repeated many times, resulting in many pairs of $(b_{\text{truth}},b_{\text{guess}})$. Then, the $\mathrm{FP}$ and $\mathrm{FN}$ for adversary’s guessing are computed by Clopper-Pearson method [10] with a confidence specification. If the adversary can make very accurate guesses and derive a lower bound higher than some claimed privacy parameter, it suggests $\mathcal{P}$ is not private as claimed.
Audit only gives a lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ of the true privacy bound $\varepsilon_{T}$, meaning that the algorithm is at least not $(\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})},\delta)$. The lower bound due to privacy audit is different from the upper bound given by theory. The limitation of privacy audit is that the result it gives should not be used as a formal privacy guarantee.
@@
Related work on privacy audit. In privacy-preserving machine learning, privacy audit mainly serves a different goal from that of certain earlier studies [52, 16, 7, 6] on detecting privacy violation in general query-answering applications. Previous work on privacy audit in machine learning mainly targets auditing the DP-SGD protocol to assess its theoretical versus practical privacy [39, 24, 23]. Additional studies [47, 37, 31, 57] concentrate on enhancing the strength of audits on DP-SGD (yielding stronger/larger-value lower bound) or improving the efficiency (incurring fewer simulation overheads). Drawing a parallel to the action of guessing whether a data point was included or not, privacy audit may also be linked to membership inference attack (MIA) [38, 44]. Still, privacy audit aims to give a privacy lower bound. There are also recent works on auditing prediction [8] and synthetic data generation [3], which differ from our auditing experiments.
II-C Private Hyper-parameter Tuning
@@
Problem Formulation. We formulate the private hyper-parameter tuning problem aligning with [30, 42]. Let $\Omega=\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{m}\}$ be a collection of DP-SGD algorithms ($m$ is chosen freely). These correspond to $m$ possible hyper-parameter configurations. We have $\mathcal{M}_{i}:\mathcal{X}\rightarrow\mathcal{Y}$ for $i\in[m]$, and all of these algorithms satisfy the same privacy parameter, e.g., they satisfy the same $(\varepsilon,\delta)$-DP guarantee.
Note that it requires that each $\mathcal{M}_{i}$ to be differentially private on its own [30, 42], meaning that indistinguishability exists between distribution $\mathcal{M}_{i}(X)$ and $\mathcal{M}_{i}(X^{\prime})$, for any $X,X^{\prime}$ being adjacent and for any $\mathcal{M}_{i}$’s hyper-parameter.
@@
Finally, it is to return an algorithm element (including its execution) of $\Omega$ such that the output of such algorithm has (approximately) the maximum score as specified by some score function $g:\mathcal{Y}\rightarrow\mathbb{R}$. The score function $g$ usually serves a utility purpose (e.g., $g$ could evaluate the validation loss on a held-out dataset). The selection must be performed in a differentially private manner. The general private selection problem corresponds to the cases where $\Omega$ contains arbitrary differentially private algorithms.
@@
Related work on private hyper-parameter tuning. Well-known algorithms like the sparse vector technique [19] and exponential mechanism [32] may potentially be leveraged to the tuning problem; however, they assume a low sensitivity in the metric defining the “best”, a condition not always applicable. Some earlier work [9] also suffers from the same issue. Papernot et al. [42] and Liu and Talwar [30] have provided generic private selection approaches circumventing such challenges. Mohapatra et al. [36] study privacy issues in adaptive hyper-parameter tuning under DP, which is different from the non-adaptive tuning problem considered in this work. There is related work [25] that integrates the solution from [42] to some larger algorithm; therefore, what we understand about the generic approach in this study naturally propagates to [25].
@@
Focus of this paper. Towards understanding how selection leaks privacy, our first focus is to formulate specific privacy audit to understand how privacy deteriorates due to selection, diverging from all previous privacy audit work on understanding privacy deteriorating due to composition. Also motivated by our empirical findings, we further improve privacy upper bound specifically for a white-box application: the hyper-parameter tuning problem, pre-conditioned on the base algorithm, is DP-SGD. There is another notable work on private hyper-parameter tuning [15] by proposing a different algorithm with different assumptions than [42] and [30] (in [42] and [30], they only require the base algorithm to be DP; in [15], they must partition the training dataset to be disjoint). Since the starting point of this work is [42] and [30], we will only focus on the same line of [42] and [30].
III Current Private Selection Protocol
III-A Current Algorithm
We start with the state-of-the-art algorithm for private selection [30, 42], shown in Algorithm 2. This generic method applies when the base algorithm is already differentially private. When each $\mathcal{M}_{i}\in\Omega$ is a DP-SGD instance, the task becomes private hyperparameter tuning. If the base algorithm $\mathcal{M}$ is $(\varepsilon,0)$-DP and $\xi$ is geometric, then Algorithm 2 is $(3\varepsilon,0)$-DP [30]. [42] improves this for pure DP by using a Truncated Negative Binomial (TNB) distribution for $\xi$ under specific parameters (see Appendix A-B). The improvement uses RDP-based analysis.
Algorithm 2 Private Selection Protocol $\mathcal{H}$ [42, 30]